OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (anonymous):

If we differentiate the distance function with respect to t, we get a function for velocity v(t). Sub t=2 into that for your answer.

OpenStudy (anonymous):

so it will be s(2)=-2-6(2)=-14 ?

OpenStudy (anonymous):

No s(t) is the distance function. Differentiate it to get a function for velocity THEN sub in t=2.

OpenStudy (anonymous):

i dont know how to diffrentiate :(

OpenStudy (anonymous):

At all or just for this question? The question mentions finding the derivative, are you behind in your class or something?

OpenStudy (anonymous):

Is this from Pre-Calculus?

OpenStudy (anonymous):

Yes it's one of my extra credit questions

OpenStudy (anonymous):

Are you familiar with this?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Go ahead and do that. :) Btw, that equation is the long method. And I can show you the shortcut if you want.

OpenStudy (anonymous):

Im listening

OpenStudy (anonymous):

I suggest that you do the equation above first since it's in your lesson. Let me know if already have the answer. :)

OpenStudy (anonymous):

wait, i am sort of confused how to plug this in

OpenStudy (anonymous):

\[f'(t)=\frac{ f(t+h)-f(t) }{ h }\] original equation: f(t)= -2-6t just add the h: f(t+h)= -2-6(t+h)

OpenStudy (anonymous):

Now substitute it to the equation.

OpenStudy (anonymous):

i plugged it in and ended up getting -4-12t

OpenStudy (anonymous):

Show me your work, so that I can tell where you got it wrong. :)

OpenStudy (anonymous):

-2-6t-6h-2-6t= -4-12t-6h= -4-12t-6h/h= -4-12t-6= -10-12t

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

Use parenthesis. :) \(\huge f'(t)=\frac{ (-2-6t-6h)-(-2-6t) }{ h }\)

OpenStudy (anonymous):

is the answer -2?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

howwww

OpenStudy (anonymous):

i just worked it out

OpenStudy (anonymous):

Just distribute the negative sign!

OpenStudy (anonymous):

oh its -6

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

THANKYOUUU

OpenStudy (anonymous):

No problem! :)

OpenStudy (anonymous):

For the shortcut. -2-6t A constant's derivative is always 0! So the derivative of -2 is 0. In -6t, notice that the exponent of the variable (in our case,t), is 1. Now, multiply the exponent to -6. Then subtract -1 from the exponent. The exponent of t is 1. So 1-1=0. Basically, the derivative of -2-6t is just -6. :)

OpenStudy (anonymous):

Wow! that is so simple :D @mathway